### v1

# Theoretical Investigation on Inductive Bias of Isolation Forest

Abstract
Isolation Forest (iForest) stands out as a widely-used unsupervised anomaly detector valued for its exceptional runtime efficiency and performance on large-scale tasks. Despite its widespread adoption, a theoretical foundation explaining iForest’s success remains unclear. This paper theoretically investigates the conditions and extent of iForest’s effectiveness by analyzing its inductive bias through the formulation of depth functions and growth processes. Since directly analyzing the depth function proves intractable due to iForest’s random splitting mechanism, we model the growth process of iForest as a random walk, enabling us to derive the expected depth function using transition probabilities. Our case studies reveal key inductive biases: iForest exhibits lower sensitivity to central anomalies while demonstrating greater parameter adaptability compared to $k$-Nearest Neighbor anomaly detectors. Our study provides theoretical understanding of the effectiveness of iForest and establishes a foundation for further theoretical exploration.
1 Introduction
Unsupervised anomaly detection is a fundamental problem in the fields of machine learning and data mining. It is widely used in many real-world tasks, including fraud detection [Fawcett and Provost, 1999, Chandola et al., 2009], network intrusion detection [Phoha, 2002], and medical diagnosis [Wong et al., 2003, Fernando et al., 2022]. The goal is to identify anomalies in the concerned dataset, which are “few and different” from the normal samples [Liu et al., 2008, 2012]. Among various types of unsupervised anomaly detection methods, iForest [Liu et al., 2012] stands out as one of the most popular choices. The basic idea is to gather a set of Isolation Trees (iTrees) trained from the observed data, and the anomaly score of a data point is defined as the average path length from the root node to the leaf node of the iTrees. The core concept of iTrees lies in uniformly randomly selecting an attribute and a split value, partitioning the data points into two subsets iteratively, and repeating this process until all data points are isolated. Although being highly heuristic, iForest is likely to be the first choice for anomaly detection. iForest outperforms other unsupervised anomaly detectors in many real-world applications in terms of accuracy, efficiency, and large-scale applicability [Liu et al., 2008, 2012, Aggarwal, 2017, Pang et al., 2019, Cook et al., 2020, Pang et al., 2022].
Despite notable progress in practical applications, the theoretical understanding of the iForest algorithm remains limited. There are some seminal works that study topics such as convergence [Siddiqui et al., 2016], PAC theory [Liu et al., 2018], and learnability [Fang et al., 2022] in the context of anomaly detection. However, these investigations are typically not specific to iForest and fail to provide a comprehensive explanation of the algorithm. Buschjäger et al. [2022] analyzed a modified version of iForest, whereas this variant deviates significantly from the original algorithm. To establish a robust theoretical foundation for the success of iForest, it is crucial to focus directly on the algorithm itself and perform comparative analyses with other anomaly detection methods.
This paper aims to advance the theoretical understanding of iForest by investigating the conditions under which it works and the extent of its efficacy. A central focus of our study is to examine the inductive bias inherent in iForest. The algorithm’s inherently stochastic growth mechanism, in which each split is selected randomly and can significantly alter the structure of an Isolation Tree, presents considerable challenges in evaluating its outcomes and identifying its inductive bias. To address this issue, we model the growth process of iTrees as a random walk, a special type of Markov chain. This perspective helps derive the expected depth function of iForest, the output of the algorithm that serves as the criterion for anomaly detection. The expected depth function enables analysis of the inductive bias and provides a valuable framework for understanding the remarkable success of iForest.
Our contributions can be summarized as follows:
  1. 1.
    We are the first to open the black box of iForest by modeling the growth process as a random walk and derive the exact depth function, which supports the further analysis of inductive bias.
  2. 2.
    We study the inductive bias of iForest by comparing it with $k$-NN through multiple case studies, concluding that iForest is less sensitive to central anomalies but more parameter-adaptive.
The rest of this paper is organized as follows. Section 2 introduces notations, settings, and concepts. Section 3 presents the novel random walk view of the growth process of iForest and the derived exact depth function. Section 4 analyzes the inductive bias of iForest through multiple case studies. Section 5 conducts some empirical studies to verify the theoretical findings. Section 6 discusses the multi-dimensional scenarios. Section 7 concludes this work. All proofs are deferred to the Appendix.
2 Preliminary
Notations. We denote by $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ if there exist constants $c_{1},c_{2}>0$ such that $f(n)\leq c_{1}g(n)$ and $f(n)\geq c_{2}g(n)$, respectively, for all sufficiently large $n\in\mathbb{N}^{+}$. Similarly, $f(n)=o(g(n))$ and $f(n)=\omega(g(n))$ if $f(n)/g(n)\to 0$ and $f(n)/g(n)\to\infty$ as $n\to\infty$, respectively. Furthermore, $f(n)=\Theta(g(n))$ if both $f(n)=O(g(n))$ and $f(n)=\Omega(g(n))$ hold simultaneously. Let $x_{i:j}$ denotes the vector $(x_{i},\dots,x_{j})$.
Settings. Let $\mathcal{X}\subset\mathbb{R}^{d}$ and $\mathbf{x}=(\mathbf{x}^{(1)},\dots,\mathbf{x}^{(d)})\in\mathcal{X}$ be the input space and a $d$-dimensional vector, respectively. We observe a set consisting of $n$ samples $D=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}\}$ drawn (allow non-i.i.d.) from an unknown distribution on $\mathcal{X}$. Dataset $D$ contains $n_{0}$ normal samples and $n_{1}$ anomalies where $n=n_{0}+n_{1}$. We have an anomaly subset $A\subsetneq D$ which is unknown and unlabeled. Anomaly detection tasks are required to find the anomaly subset in the given dataset $D$.
In unsupervised anomaly detection tasks, however, ground-truth labels are unavailable, making the task inherently subjective. Different individuals may consider different anomaly subsets, depending on one’s preferences. There are no loss functions to optimize, and understanding the inductive biases of various anomaly detection algorithms is crucial, rather than focusing solely on their absolute performance. In this study, we compare two widely used anomaly detectors: iForest and $k$-NN.
Algorithm 1 BuildTree$(D)$
Input: A dataset $D=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots\}$
Output: An Isolation Tree $T$
1:if $|D|\leq 1$ then
2:     return Leaf
3:$j\leftarrow$ uniform random in $\{j\mid\#\{\mathbf{x}^{(j)}\}>1\}$
4:$s\leftarrow$ uniform random in $[\min\mathbf{x}^{(j)},\max\mathbf{x}^{(j)}]$
5:$D_{\text{left}}\leftarrow\{\mathbf{x}\in D\mid\mathbf{x}^{(j)}\leq s\}$
6:$D_{\text{right}}\leftarrow\{\mathbf{x}\in D\mid\mathbf{x}^{(j)}>s\}$
[…]
8:$\quad\quad\quad\quad~{}\text{SplitPoint}\leftarrow s$
9:$\quad\quad\quad\quad~{}\text{Left}\leftarrow\text{BuildTree}(D_{\text{left}})$
[…]
11:return Node
Algorithm 2 Depth $(\mathbf{x},T)$
Input: A sample $\mathbf{x}$ and an Isolation Tree $T$
Output: The depth $h(\mathbf{x},T)$
1:if $T$ is a Leaf then
2:     return 00
3:$j\leftarrow T.\text{SplitAtt},s\leftarrow T.\text{SplitPoint}$
4:if $\mathbf{x}^{(j)}\leq s$ then
5:     return $1+\text{Depth}(\mathbf{x},T.\text{Left})$
6:else
7:     return $1+\text{Depth}(\mathbf{x},T.\text{Right})$
Algorithm 3 Isolation Forest
Input: A dataset $D$ and the tree number $M$
Output: A score function
1:for $m=1,\dots,M$ do
2:     $h_{m}\leftarrow\textrm{Depth}(\cdot,\textrm{BuildTree}(D))$
3:return $M^{-1}\sum_{m=1}^{M}h_{m}$
iForest. The concept of iForest [Liu et al., 2008] has exceeded the scope of anomaly detection and become a learning framework for diverse machine learning tasks, including density estimation [Ting et al., 2021], time-series analysis [Ting et al., 2022], and so on. There are also many variants of iForest, such as SCiForest [Liu et al., 2010], LSHiForest [Zhang et al., 2017], Deep Isolation Forest [Xu et al., 2023], OptIForest [Xiang et al., 2023], etc. Note that theoretically understanding the large family of isolation-based methods is a long way to go. This paper focuses on the original iForest [Liu et al., 2008], which is the basis of understanding the idea of isolation as well as the large algorithm family.
iForest works under the belief that anomalies are easier to isolate than normal data points if we uniformly randomly partition the whole feature space. iForest constructs a collection of iTrees independently, and the anomaly score is negatively correlated with the average depth, the path length from the root node to the leaf node. The procedure of iForest is detailed in Algorithms 1-3.
For convenience, we introduce two simplifications: relaxing the maximum depth of iTrees and directly outputting the average depth [Liu et al., 2008] instead of computing the anomaly score. About the former simplification, constraining the maximum depth will not alter the relative order of isolation depths for different data points and is usually applied to detect multiple anomaly candidates [Liu et al., 2008], which is not the focus of this work. As for the latter simplification, note that the anomaly score introduced in Liu et al. [2008]
$s(\mathbf{x},n)=2^{-\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]/c(n)}$
constitutes a monotonically decreasing transformation of the average depth. This shifts the focus from analyzing maximum anomaly scores to minimum depths without altering the concerned problem.
$k$-NN anomaly detectors. Nearest neighbor-based methods can be roughly separated into two categories: density-based methods that assume the anomalies are in low-density regions and distance-based methods that assume the anomalies are far from normal points. Among the former, the most popular ones include Local Outlier Factor [Breunig et al., 2000], and its variants [Tang et al., 2002, Papadimitriou et al., 2003, Fan et al., 2006]. Among the latter, the most typical ones define the anomaly score as the distance to the $k$-th nearest neighbors of a point [Byers and Raftery, 1998, Guttormsson et al., 1999]. Here, we consider the most standard $k$-NN, which defines the anomaly score as the average distance to the $k$ nearest neighbors. Formally, the scoring function is defined by
$h_{knn}(\mathbf{x};D)\triangleq\frac{1}{k}\sum_{\mathbf{x}^{\prime}\in\mathcal{N}_{k}(\mathbf{x})}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\ ,$
where $\mathcal{N}_{k}(\mathbf{x})$ is the set of the $k$ nearest neighbors of $\mathbf{x}$ in $D$. The choice of the $L_{1}$-norm is representative because all norms in finite-dimensional spaces are equivalent. The reason we select the $L_{1}$-norm is that each iTree partitions the entire feature space into multiple hyper-rectangles. This characteristic aligns more closely with the $L_{1}$-norm than the more commonly used $L_{2}$-norm.
Here, we employ the $k$-NN algorithm rather than density-based approaches. It becomes apparent that density-based algorithms employing the rectangle kernel $K(u)=2^{-d}~{}\mathbb{I}(\|u\|_{1}\leq 1)$ tend to be equivalent to $k$-NN. This equivalence arises due to the strong correlation between the number of points in the neighborhood of one point and the distance to its nearest neighbors.
3 Expected depth function of iForest
In this section, we conduct a mathematical analysis of the depth function of iForest. The depth function, the scoring criterion output by the iForest, serves as a basis for the analysis. Note that it is challenging to analyze the depth function of iForest due to its unpredictability in the growth process. We tackle this challenge by modeling the growth of iTree as a random walk and derive the closed-form expression of depth functions.
3.1 The random walk model for iTrees
We denote the average depth of a point $\mathbf{x}$ in an iTree as $h(\mathbf{x};D,\mathbf{\Theta})$, where $\mathbf{\Theta}$ consists of all the randomness from the construction of the iTree, including the random subsampling, random attribute selections, and random split values. To begin with, we introduce the following conclusion.
Proposition 3.1 (Concentration of iForest).
For any fixed dataset $D$ and any $\mathbf{x}\in\mathcal{X}$, we have
$\Pr\Bigg{[}\Bigg{|}\frac{1}{M}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})-\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]\Bigg{|}\geq\epsilon\Bigg{]}\leq 2\exp\left(-2\epsilon^{2}M/n^{2}\right)\ ,$
where $h(\mathbf{x};D,\mathbf{\Theta}_{m})$ is the depth of $\mathbf{x}$ in the $m$-th tree.
Proposition 3.1 shows that the empirical mean of the depth function $M^{-1}\sum_{m=1}^{M}h(\mathbf{x};D,\mathbf{\Theta}_{m})$ will with high probability converges to its expectation $\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};D,\mathbf{\Theta})]$ as the number of trees $M$ increases. Following the proposition, it suffices to analyze the expected depth function instead of the empirical mean. However, the generating process of iTrees is quite unpredictable, coming from two aspects: Every split is chosen randomly but depends on the given dataset; a different split can lead to a completely different tree. Thus, it is still a great challenge to analyze the expected depth function.
Next, we demonstrate that the process of generating an iTree can be modeled as a random walk. For simplicity, we first consider the case of $d=1$, i.e., $x_{i}\in\mathcal{X}=\mathbb{R}$. Without loss of generality, we assume that $D=\{x_{1},x_{2},\dots,x_{n}\}$ is sorted in ascending order, i.e., $x_{1}<x_{2}<\dots<x_{n}$. Suppose we care about the depth of $x_{i}$, let $x_{\ell_{t}}$ and $x_{r_{t}}$ denote the smallest and the largest point in the tree node containing $x_{i}$ at $t$, respectively. The following theorem shows that $\mathbf{s}_{t}=(x_{\ell_{t}},x_{r_{t}})$ is a random walk.
Theorem 3.2 (Random Walk Model for iTree).
For any $x_{\ell},x_{r},x_{\ell^{\prime}},x_{r^{\prime}}\in D$, random process $\mathbf{s}_{t}(x_{i})\triangleq(x_{\ell_{t}},x_{r_{t}})$ is a random walk with transition probability
$\Pr[\mathbf{s}_{t+1}=(x_{\ell^{\prime}},x_{r^{\prime}})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})]=\left\{\begin{aligned} &\frac{x_{\ell^{\prime}}-x_{\ell^{\prime}-1}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}>\ell,r^{\prime}=r\ ,\\ &\frac{x_{r^{\prime}+1}-x_{r^{\prime}}}{x_{r}-x_{\ell}}\ ,&\text{if }\ell^{\prime}=\ell,r^{\prime}<r\ ,\\ &0\ ,&\text{otherwise}\ .\end{aligned}\right.$ (1)
Figure 1: An example of the random walk model for an iTree. The points colored in black, blue, and red indicate the initial, intermediate, and absorbing states, respectively.
(a) Marginal single anomaly.
(b) Central single anomaly.
(c) Marginal multiple anomalies.
Figure 2: Three types of concerned anomalies in the case studies. Red and blue points represent anomalies and normal points, respectively.
Here, we omit the absorbing state $(x_{i},x_{i})$ for simplicity, which will with probability one keeping its state. Figure 3.1 shows an example of the random walk model for an iTree, where the original dataset is $D=\{x_{1},x_{2},\dots,x_{7}\}$, of which the relative positions are shown in the two axes. Every state of the random walk can be represented by a point in the two-dimensional coordinate system. The initial state of the random walk is $\mathbf{s}_{0}=(x_{1},x_{7})$, which is represented by the black points. At each step, it is allowed to move to the right or up, keeping at the left and below the absorbing point $(x_{4},x_{4})$. For example, the random walk reaches the state $(x_{3},x_{6})$ from $(x_{1},x_{7})$ by moving rightward and then upward, which is represented by the blue points. The probability of moving how large the step size is proportional to the distance between two points, as defined in Eq. (1). The random walk keeps running until it reaches the absorbing point $(x_{4},x_{4})$, which is represented by the red points.
Though we only show the random walk model when $d=1$, it still holds for any $d>1$. This is because the candidate split attribute and point only depend on the current candidate points, no matter how the current state is reached, i.e., the growth of iTrees verifies Markov property.
3.2 Expected depth of a data point
With the random walk model at hand, we can now analyze the expectation of the depth function. We begin with a simpler notation of the expected depth function as follows
$\bar{h}(x_{i};x_{1},\dots,x_{n})\triangleq\mathbb{E}_{\mathbf{\Theta}}[h(x_{i};x_{1},\dots,x_{n},\mathbf{\Theta})]\ ,$
where $x_{i}\in D$. We then have the following lemma.
Lemma 3.3.
For any given dataset $D$ with sample size $n\geq 3$ and $x_{i}\in D$ for $1<i<n$, we have
$\bar{h}\left(x_{i};x_{1},\dots,x_{n}\right)=\bar{h}\left(x_{i};x_{1},\dots,x_{i}\right)+\bar{h}\left(x_{i};x_{i},\dots,x_{n}\right)\ .$
Lemma 3.3 shows that the expected depth of a data point $x_{i}$ can be decomposed into two parts, of which the concerned data point is the rightmost and the leftmost, respectively. From the random walk perspective, the anticipated number of steps required to reach the absorbing point $(x_{i},x_{i})$ equals the sum of the steps taken in the rightward direction and the steps taken upward. For example, although opting to move rightward can increase the probability of moving upward in the next step, the probability distribution of the step size of moving, conditioning on moving upward, remains unchanged; thus, the expected total steps remain unchanged, too.
From the above analysis, it suffices to analyze the expected depth of the rightmost point $x_{i}$ in $\{x_{1},\dots,x_{i}\}$, as the leftmost point can be analyzed similarly. By the transition rule of random walks, the depth of point $x_{i}$ has a probability cumulative distribution function as follows.
$\Pr_{\mathbf{\Theta}}[h(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta})\leq\xi]=[1,0,\dots,0]\cdot\mathbf{P}^{\xi}\cdot[0,0,\dots,1]^{\top}\ ,$
where $\mathbf{P}=[p_{jk}]_{i\times i}$ is the transition matrix with
$p_{jk}=\Pr[\mathbf{s}_{t+1}=(x_{k},x_{i})\mid\mathbf{s}_{t}=(x_{j},x_{i})]$
and $\xi\in\mathbb{N}^{+}$. Thus, the expected depth of $x_{i}$ equals
$\bar{h}(x_{i})=\sum_{\xi=1}^{i}\Pr_{\mathbf{\Theta}}[h(x_{i};\mathbf{\Theta})\leq\xi]\ ,$
where $\bar{h}(x_{i};x_{1},\dots,x_{i})$ and $h(x_{i};x_{1},\dots,x_{i})$ are abbreviated as $\bar{h}(x_{i})$ and $h(x_{i})$, respectively. By some algebraic manipulations, we have the closed-form expression of the expected depth below.
Lemma 3.4.
For any given dataset $D=\{x_{1},\dots,x_{i}\}$ with sample size $i>1$, we have
$\bar{h}(x_{i};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{i}-x_{j-1}}\quad\textrm{and}\quad\bar{h}(x_{1};x_{1},\dots,x_{i})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{j}-x_{1}}\ .$
Lemma 3.4 presents the closed-form expression of the expected depths of the rightmost or leftmost point. Based on a straightforward combination of Lemma 3.3 and Lemma 3.4, we have the expected depth of any $x_{i}$ in $D$ in the following theorem.
Theorem 3.5.
For any given dataset $D$ with sample size $n>2$ and $x_{i}\in D$, we have
$\bar{h}(x_{i};x_{1},\dots,x_{n})=\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{i}-x_{j-1}}+\sum_{j=i+1}^{n}\frac{x_{j}-x_{j-1}}{x_{j}-x_{i}}\ .$
Theorem 3.5 presents the closed-form expression for the depth function of any data point within a given dataset. This result represents the first theoretical characterization of the decision function generated by iForest, providing a foundational basis for further analysis of the inductive bias.
For any unseen test data points, the depth function is the linear interpolation of the depths of the two adjacent data points, which is beyond the scope of this paper. For anyone interested, we refer the reader to Theorem A.3, where we provide the detailed conclusion and proof.
4 Case studies for inductive bias
Based on the closed-form expression of expected depth functions, we now conduct case studies to compare the inductive bias of iForest with $k$-NN. Before that, we define the density metrics.
Definition 4.1 (Density Metrics).
Let $U\triangleq\max_{i\geq 2}~{}\lvert x_{i+1}-x_{i}\rvert$ and $L\triangleq\min_{i\geq 2}~{}\lvert x_{i+1}-x_{i}\rvert$ be the maximum and minimum distances between adjacent points, respectively. Define two density metrics:
  • The density factor $\kappa$ measures the ratio of $U$ to $L$: $\kappa=U/L$.
  • The density difference $\delta$ measures the difference between $U$ and $L$: $\delta=U-L$.
These metrics evaluate how uniformly data points are distributed relative to their neighbors. Intuitively, a dataset composed entirely of normal points is expected to exhibit small values for both metrics; otherwise, some points would be excessively distant from their neighbors. In the case studies below, we demonstrate that these density metrics play a critical role in shaping the inductive bias of iForest.
Before proceeding, we introduce the following assumption.
Assumption 4.2.
Assume that density factor $\kappa\geq\sqrt{n+3}$.
Assumption 4.2 is frequently observed in practice. To justify its reasonability, we check all dimensions of binary classification datasets from the OpenML benchmark [Vanschoren et al., 2014, Bischl et al., 2021]. The results are summarized in Table 1. “Successful” indicates that the density factor exceeds $\sqrt{n+3}$. “Valid” indicates the nonexistence of repeated values in the dimension. Notably, the theoretical probability of repeated values is zero. Thus, we omit this case for the convenience of analysis. Our analysis demonstrates that the density factor exceeds $\sqrt{n+3}$ in almost all dimensions across all datasets, highlighting the mildness of Assumption 4.2. A theoretical study of Assumption 4.2 is also conducted, with detailed results available in Appendix B.
Table 1: Number of features in OpenML benchmark that verify Assumption 4.2.
Successful Valid Total
930,738 930,751 933,440
4.1 Marginal single anomaly
The first case study is about the marginal single anomaly, an example of which is shown in Figure 2 (a), where the anomaly is located at the leftmost and the normal points are distributed at the right side. Marginal single anomalies may be the most common ones in practice, which happens when there is some noise in normal data drawn from continuous and bounded support. Note that although this model looks simple, it is the basic building block of more complex models.
For the marginal single anomaly, we have the following theorem for iForest.
Theorem 4.3.
Suppose that $x_{2}-x_{1}>U\cdot\kappa$. For any dataset $D$ with $x_{2:n}$ verifying $\kappa$-dense, it holds that $\bar{h}(x_{1};x_{1:n})<\bar{h}(x_{j};x_{1:n})$ for all $j>1$.
Theorem 4.3 shows that the expected depth of the marginal single anomaly is the smallest among all the data points when $x_{2}-x_{1}>U\cdot\kappa$. Hence, the marginal single anomaly is the most likely to be the anomaly output by iForest. For the rightmost marginal single anomaly Theorem 4.3 holds similarly. Theorem 4.3 shows the sufficiency of $x_{2}-x_{1}>U\cdot\kappa$. The following theorem shows the necessity.
Theorem 4.4.
Suppose that $U<x_{2}-x_{1}\leq U\cdot\kappa$, and $x_{2:n}$ verifies $\kappa$-dense. For any $n>4$, there exists an assignment to $x_{1},\dots,x_{n}$ such that $\bar{h}(x_{1};x_{1:n})\geq\bar{h}(x_{j_{0}};x_{1:n})$ holds for some $j_{0}>1$.
Theorem 4.4 shows that if $x_{2}-x_{1}>U\cdot\kappa$ is violated, even though $x_{2}-x_{1}$ is the largest distance between two consecutive points, it remains possible for $x_{1}$ not to be the shallowest point, or in other words, iForest may fail to detect the marginal single anomaly.
For $k$-NN, we have the following conclusion.
Theorem 4.5.
Suppose that $x_{2:n}$ verifies $\delta$-dense. The sufficient and necessary condition for $k$-NN to detect the marginal single anomaly is $x_{2}-x_{1}>U+(k-1)\delta/2$.
Theorem 4.5 reveals that the threshold for $k$-NN depends on the choice of $k$; when $k$ is too large, the marginal single anomaly is likely to be missed, while choosing a small $k$ may ignore some global information. In summary, Theorems 4.3-4.5 imply that isolation-based anomaly detection is more adaptive to marginal single anomalies than $k$-NN.
4.2 Central single anomaly
Central single anomaly is another common type of anomaly, as shown in Figure 2 (b). The anomaly is located at the center of the data points and the normal points are distributed at both the left and right sides. This models the scenario when the data distribution consists of multiple clusters, and the anomaly is likely to arise in the gap between different clusters.
For iForest, we have the following theorem.
Theorem 4.6.
Let $n_{0}$ be an even number for convenience. Suppose that both $x_{1:n_{0}/2}$ and $x_{(n_{0}/2+2):n}$ verify $\kappa$-dense. Then, the sufficient and necessary condition for iForest to detect $x_{n_{0}/2+1}$ as an anomaly is that $\min\{x_{n_{0}/2+1}-x_{n_{0}/2},x_{n_{0}/2+2}-x_{n_{0}/2+1}\}>\Theta(\sqrt{n_{0}\kappa})$.
Theorem 4.6 shows the sufficiency and necessity for iForest detecting the central single anomaly. Here, the anomaly is “different” enough from the normal points to a degree of order $\Theta(\sqrt{n_{0}\kappa})$, which is a relatively large quantity, especially when $n_{0}$ is large. This may be because iForest is more likely to capture global information for data distribution.
For $k$-NN, we have the following theorem.
Theorem 4.7.
Let $n_{0}$ be an even number for convenience. Suppose that both $x_{1:n_{0}/2}$ and $x_{(n_{0}/2+2):n}$ verify $\delta$-dense. The sufficient and necessary condition for $k$-NN to detect the central single anomaly is that $\min\{x_{n_{0}/2+1}-x_{n_{0}/2},x_{n_{0}/2+2}-x_{n_{0}/2+1}\}>\Theta(k\delta)$.
Theorem 4.7 shows that the threshold for $k$-NN also relies on the selection of $k$, which is an algorithm-dependent parameter, while the threshold for iForest relies only on the data distributions, which is problem-dependent. We conclude that iForest is more careful about detecting central single anomalies but is also more parameter-adaptive than $k$-NN.
4.3 Marginal clustered anomalies
In this subsection, we consider the case of marginal clustered anomalies. The motivation is that anomaly events may occur multiple times, as shown in Figure 2 (c). A group of anomalies is located at the leftmost of the data points, and the normal points are distributed on the right side. Marginal clustered anomalies characterize some unimodal data with multiple clustered noisy points.
For marginal clustered anomalies, we have the following conclusion for iForest.
Theorem 4.8.
Let $n_{1}$ be an odd number for convenience. Suppose $n_{1}=o(n_{0})$. For any dataset $D$, of which $x_{1:n_{1}}$ and $x_{(n_{1}+1):(n_{1}+n_{0})}$ verify $\kappa$-dense, the sufficient and necessary condition for iForest to detect all the marginal clustered anomalies is that $x_{n_{1}+1}-x_{n_{1}}>\Theta(n_{1}^{2}\kappa)$.
Theorem 4.8 shows that iForest can detect all the clustered anomalies when normal points are far enough from anomalies, while it requires a quite large threshold of order $\Theta(n_{1}^{2}\kappa)$ when $n_{1}$ is large.
For $k$-NN, we have the following theorem.
Theorem 4.9.
Suppose $\omega(n_{1})\leq k\leq o(n_{0})$. $x_{1:n_{1}}$ and $x_{(n_{1}+1):(n_{1}+n_{0})}$ verify $\delta$-dense. The sufficient and necessary condition for $k$-NN to detect all marginal clustered anomalies is $x_{n_{1}+1}-x_{n_{1}}>\Theta(k\delta)$.
Theorem 4.9 shows that the threshold for $k$-NN is also algorithm-dependent, yielding properties similar to those in Theorems 4.5 and 4.7. The selection of $k$ holds considerable importance for the $k$-NN. For instance, when $k$ is too small, $k$-NN may wrongly label the clustered anomalies as normal points, leading to undesirable missed detections and false positives. In contrast, iForest keeps exhibiting adaptability without any hyperparameter tuning. Note that the parameters of iForest are often set to trade off between the approximation performance and the computational cost, while the parameters of $k$-NN are set to trade off between the false positives and the false negatives.
Table 2 summarizes three cases of anomalies in this section. It is observed that iForest acts to be more cautious in reporting an anomaly as the threshold is relatively large but more parameter-adaptive to different types of anomalies than $k$-NN as the thresholds of iForest are only problem-dependent.
Table 2: Decision thresholds of iForest (IF) and $k$-NN for different types of anomalies.
Anomaly type IF $k$-NN
Marginal single $U\cdot\kappa$ $\Theta(k\delta)$
Central single $\Theta(\sqrt{n_{0}\kappa})$ $\Theta(k\delta)$
Marginal clustered $\Theta(n_{1}^{2}\kappa)$ $\Theta(k\delta)$
(a) Synthetic datasets
(b) HTTP dataset
(c) SMTP dataset
Figure 3: Mean-square errors of the real depths learned by iForest to the theoretical expected depths about numbers of trees. The shaded regions represent 95% confidence regions over multiple runs.
(a) Example 1
(b) Example 2
(c) Example 3
Figure 4: Visualization of the expected depths for points generated from a uniform distribution. Circles and bars represent the point positions and the expected depths, respectively.
5 Empirical studies
In this section, we conduct experiments to validate our theoretical findings. These include examining the convergence of the empirical depth and when iForest predicts a point as an anomaly.
5.1 Convergence of the empirical depth
This experiment is conducted using both synthetic and real-world datasets. We begin by independently sampling 100 points from three distributions: the standard normal distribution $\mathcal{N}(0,1)$, the uniform distribution $U[0,1]$, and the exponential distribution $\textrm{Exp}(1)$. Additionally, we examine the HTTP and SMTP datasets [UCI KDD Archive, 1999], which are widely used anomaly detection datasets. Focusing on the one-dimensional case, we separately select each feature from all datasets to demonstrate the convergence of the empirical depth. For each dataset, we evaluate the mean squared errors of the depths learned by iForest to the theoretical expected depths with 100 subsamples as the number of iTrees ranges from 100 to 1000. Each setup is repeated 10 times independently.
Figure 3 shows the curves of the errors about the number of iTrees. The error decreases rapidly as the number of trees increases, which implies that the depth of every point may eventually converge to the theoretically expected depth. This verifies the concentration property of the empirical depth and the correctness of the expected depth function shown in Prososition 3.1 and Theorem 3.5, respectively.
5.2 Anomaly detection of iForest
We generate multiple datasets, each of which contains 20 points drawn from a uniform distribution independently. We compute the expected depth of each point to help understand the behavior of iForest. The results are presented in Figure 4, where we label the points with a small depth considered as anomalies and a larger depth considered as normal points in red and blue colors, respectively.
From Figure 4, three key observations can be made: (1) Marginal single and clustered anomalies are correctly labeled, even though some are not significantly distant from normal points. This aligns with Theorems 4.3 and 4.8, stating that marginal anomalies are more easily isolated. (2) In Figure 4 (a), only the central anomaly is detected. This is consistent with Theorem 4.6, which indicates that iForest can identify central anomalies only when they are sufficiently distant from normal data. (3) In Figures 4 (b) and (c), $k$-NN may fail to detect the rightmost clustered anomaly when $k$ is set too small, whereas iForest successfully identifies all marginal clustered anomalies. This supports Theorems 4.84.9, demonstrating that iForest is more adaptive to parameter settings than $k$-NN.
6 Discussions
The above analysis focuses on one-dimensional cases. In this section, we extend to multi-dimensional cases, including reducible and irreducible cases.
(a) Reducible case 1
(b) Reducible case 2
(c) Irreducible case
Figure 5: The expected depth function in the reducible and irreducible cases.
Reducible case. We say that the anomalies are reducible if they remain anomalous when projected on any dimension. For example, Figures 5 (a) and (b) can be reduced to marginal and central anomalies when projected on both dimensions, respectively. In this case, we can estimate the expected depth by the average depth projected on each dimension, defined as
$\bar{h}(\mathbf{x};\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx\frac{1}{d}\sum_{j=1}^{d}\bar{h}(\mathbf{x}^{(j)};\mathbf{x}_{1}^{(j)},\dots,\mathbf{x}_{n}^{(j)})\ .$
Irreducible case. If the anomalies are irreducible, we can first apply some feature mapping $\phi:\mathbb{R}^{d}\to\mathbb{R}$ and estimate the expected depth by the depth of $\phi(\mathbf{x})$, defined as
$\bar{h}(\mathbf{x};\mathbf{x}_{1},\dots,\mathbf{x}_{n})\approx\bar{h}(\phi(\mathbf{x});\phi(\mathbf{x}_{1}),\dots,\phi(\mathbf{x}_{n}))\ .$
If the feature mapping $\phi$ can suitably model the data distribution, one can obtain an ideal estimate of the expected depth. For example, a radial basis function (RBF) kernel mapping perfectly makes the data highly separable, as shown in Figure 5 (c). This also reduces to marginal single anomalies.
Through direct reduction or kernel mapping, we can reduce the multi-dimensional anomaly detection to one-dimensional cases and further extend our one-dimensional analysis. Hence, all the case studies and conclusions in Section 4 can be applied to the multi-dimensional cases.
7 Conclusions
We theoretically investigated the inductive bias of iForest. Specifically, we introduced a random walk model for the growth process of iForest. This model facilitates the derivation of the closed-form expected depth function of iForest, allowing for a thorough comparison with the $k$-NN anomaly detector. Through case studies, we demonstrated that iForest exhibits less sensitivity in identifying central anomalies and greater parameter adaptability compared to $k$-NN. Numerical experiments verified our theoretical findings, including the concentration of iForest and the case studies.
Appendix A Proofs in Section 3
Here, we provide the proofs of results in Sections 3.
A.1 Proof of Proposition 3.1
We begin with the following technical lemma.
Lemma A.1 (Hoeffding’s inequality [Hoeffding, 1963]).
Let $X_{1},\dots,X_{n}$ be independent random variables such that $a_{i}\leq X_{i}\leq b_{i}$ for all $i\in[n]$. Consider the sum of these random variables $S_{n}=\frac{1}{n}\sum_{i=1}^{n}X_{i}$. Then, for all $\epsilon>0$, we have
$\Pr\left[\left|S-\mathbb{E}[S]\right|\geq\epsilon\right]\leq 2\exp\left(-2\epsilon^{2}/\sum_{i=1}^{n}(b_{i}-a_{i})^{2}\right)\ .$
Then we prove Proposition 3.1.
Proof   We first denote $h^{\prime}_{m}=M^{-1}h_{m}(\mathbf{x};\mathbf{\Theta}_{m})$. From the growth process of Isolation Trees, we have that $h^{\prime}_{m}$ verifies i.i.d. condition, $h^{\prime}_{m}\in(0,n/M)$, and
$\mathbb{E}_{\mathbf{\Theta}}[h^{\prime}_{m}]=\mathbb{E}_{\mathbf{\Theta}}[h(\mathbf{x};\mathbf{\Theta})],~{}\forall 1\leq m\leq M\ .$
By applying Hoeffding’s inequality (Lemma A.1) to $h^{\prime}_{m}$, we have
$\Pr\left[\left|\sum_{m=1}^{M}h^{\prime}_{m}-\mathbb{E}\left[\sum_{m=1}^{M}h^{\prime}_{m}\right]\right|\geq\epsilon\right]$ $\leq 2\exp\left(-2\epsilon^{2}/(M(n/M)^{2})\right)$
$=2\exp\left(-2\epsilon^{2}M/n^{2}\right)\ .$
[…]
A.2 Proof of Theorem 3.2
Markov property. The Markov property holds directly from the growth process of Isolation Trees.
Transition probability. We now verify the transition probability. Recall that the current state, the next state, and the absorbed state are denoted by $(x_{\ell},x_{r})$ and $(x_{\ell^{\prime}},x_{r^{\prime}})$, $(x_{i},x_{i})$, respectively, where $x_{\ell},x_{\ell^{\prime}},x_{r},x_{r^{\prime}},x_{i}\in D$, $x_{\ell}\leq x_{i}\leq x_{r}$, and $x_{\ell^{\prime}}\leq x_{i}\leq x_{r^{\prime}}$. In every step, Isolation Tree chooses the split value in the interval $(x_{\ell},x_{i})$ or interval $(x_{i},x_{r})$, and thus either $x_{\ell^{\prime}}=x_{\ell}$ or $x_{r^{\prime}}=x_{r}$ holds, implying that
$\Pr\left[\mathbf{s}_{t+1}=(x_{\ell^{\prime}},x_{r^{\prime}})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})\right]=0\ ,$
where $x_{\ell^{\prime}},x_{r^{\prime}}\in\{(x,y)\mid(x\neq x_{\ell}\wedge y\neq x_{r})\vee(x=x_{\ell}\wedge y\geq x_{r})\vee(y=x_{r}\wedge x\leq x_{\ell})\}$.
We now consider the case of $x_{\ell^{\prime}}>x_{\ell}$ and $x_{r^{\prime}}=x_{r}$. The state change from $(x_{\ell},x_{r})$ to $(x_{\ell^{\prime}},x_{r})$ happens when the split point is chosen in the interval $(x_{\ell^{\prime}-1},x_{\ell^{\prime}})$. Hence, the transition probability equals the probability of choosing the split point in the interval $(x_{\ell^{\prime}-1},x_{\ell^{\prime}})$, which equals
$\Pr\left[\mathbf{s}_{t+1}=(x_{\ell^{\prime}},x_{r})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})\right]=\frac{x_{\ell^{\prime}-1}-x_{\ell^{\prime}}}{x_{r}-x_{\ell}},~{}\forall(x_{\ell}<x_{\ell^{\prime}}\leq x_{i})\ .$
Similarly, for $x_{\ell^{\prime}}=x_{\ell}\textrm{ and }x_{r^{\prime}}<x_{r}$, we have
$\Pr\left[\mathbf{s}_{t+1}=(x_{\ell},x_{r^{\prime}})\mid\mathbf{s}_{t}=(x_{\ell},x_{r})\right]=\frac{x_{r^{\prime}+1}-x_{r^{\prime}}}{x_{r}-x_{\ell}},~{}\forall(x_{i}\leq x_{r^{\prime}}<x_{r})\ .$
[…]
A.3 Proof of Lemma 3.4
We will prove Lemma 3.4 first before proving Lemma 3.3.
Proof   By the definition of $\bar{h}$, it is equivalent to show that
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
Let $\mathbf{\Theta}=(\Theta_{1},\dots,\Theta_{\tau})$ be the vector of random variables representing the split at each step, where $\tau$ denotes the depth of point $x_{i}$. Note that $\tau$ can be indeterminate, as the depth is a random variable. Let $\Omega$ be the event that the first split is between $x^{\prime}$ and $x_{1}$. Then, we can expand the expectation as follows:
$\mathbb{E}_{\mathbf{\Theta}}\Big{[}h\big{(}x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\big{)}\Big{]}$ $=\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[\mathbb{E}_{\Theta_{1}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\right]$
$=\Pr_{\Theta_{1}}\left[\Theta_{1}\in\Omega\right]~{}\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]$
$\quad\Pr_{\Theta_{1}}\left[\Theta_{1}\notin\Omega\right]~{}\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]$
$=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]\right)+$
$\quad\frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]\right)\ .$
For $\Theta_{1}\in\Omega$, the tree growth process after the first split is the same as the case where $x^{\prime}$ is not added
$\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]=\mathbb{E}_{\mathbf{\Theta}^{\prime}_{1:\tau}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
For $\Theta_{1}\notin\Omega$, in fact, the tree growth process is the same as the case where $x^{\prime}$ is added, i.e.,
$1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]=\mathbb{E}_{\mathbf{\Theta}^{\prime}_{1:\tau}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
Therefore, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]$ $=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\in\Omega\right]\right)+$
$\quad\frac{x_{i}-x_{1}}{x_{i}-x^{\prime}}~{}\left(1+\mathbb{E}_{\mathbf{\Theta}_{2:\tau}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\mid\Theta_{1}\notin\Omega\right]\right)$
$=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]\ .$
[…]
A.4 Proof of Lemma 3.3
The verify We begin with the following lemma.
Lemma A.2.
For any given points $x_{1}<x_{2}<\dots<x_{i}$ and $x^{\prime}<x_{1}$, we have for all $1\leq i\leq n$
$\bar{h}\left(x_{i};x^{\prime},x_{1},\dots,x_{n}\right)=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\bar{h}\left(x_{i};x_{1},\dots,x_{n}\right)\ .$
Proof   Assume that $n>i$, otherwise it recovers Lemma 3.4. By the definition of $\bar{h}$, we are to show that
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]\ .$ (2)
Applying mathematical induction on $n$. For $n=i+1$, we have the base case holds as follows:
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i+1},\mathbf{\Theta}\right)\right]$ $=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]+1$
$=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i},\mathbf{\Theta}\right)\right]+1$
$=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{i+1},\mathbf{\Theta}\right)\right]\ ,$
where the first equality follows from the fact that the new point $x_{i+1}$ increases the depth of $x_{i}$ by 1, and the second equality follows from Lemma 3.4.
Assuming that Eq. (2) holds for all $n<k\ (k>i+1)$, we then consider $n=k$. There are three cases: If the first split occurs in $[x^{\prime},x_{1}]$, the expected depth of $x_{i}$ increases by 1; If the first split occurs in $[x_{1},x_{i}]$, the expected depth of $x_{i}$ remains unchanged; otherwise, a split in $[x^{\prime},x_{i}]$ will eventually occur, and the expected depth of $x_{i}$ increases by about $\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}$. Formally, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]$ $=\frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}(1+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right])$
$\quad\quad+\sum_{j=2}^{i}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]$
$\quad\quad+\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]\ .$
By mathematical induction, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]=\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]\ ,$
which implies that
$\sum_{j=i+1}^{k}\frac{x_{j}-x_{j-1}}{x_{k}-x^{\prime}}\mathbb{E}_{\mathbf{\Theta}}$ $\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{j-1},\mathbf{\Theta}_{2:\tau}\right)\right]$
$=\frac{x_{k}-x_{i}}{x_{k}-x^{\prime}}\cdot\left(\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]\right)\ .$
Therefore, we conclude that
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x^{\prime},x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]=$ $\ \frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}+\frac{x_{k}-x_{i}}{x_{k}-x^{\prime}}\cdot\frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+$
$\ \left(\frac{x_{1}-x^{\prime}}{x_{k}-x^{\prime}}+\frac{x_{i}-x_{1}}{x_{k}-x^{\prime}}+\frac{x_{k}-x_{i}}{x_{k}-x^{\prime}}\right)\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]$
$=$ $\ \frac{x_{1}-x^{\prime}}{x_{i}-x^{\prime}}+\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{i};x_{1},\dots,x_{k},\mathbf{\Theta}\right)\right]\ .$
[…]
Then, we prove Lemma 3.3.
Proof   By Lemma A.2, we have
$\bar{h}\left(x_{i};x_{1},\dots,x_{n}\right)$ $=\bar{h}\left(x_{i};x_{1},\dots,x_{n-1}\right)+\frac{x_{n}-x_{n-1}}{x_{n}-x_{i}}$
$=\bar{h}\left(x_{i};x_{1},\dots,x_{n-2}\right)+\frac{x_{n-1}-x_{n-2}}{x_{n-1}-x_{i}}+\frac{x_{n}-x_{n-1}}{x_{n}-x_{i}}$
$=\bar{h}\left(x_{i};x_{1},\dots,x_{i}\right)+\frac{x_{i+1}-x_{i}}{x_{i+1}-x_{i}}+\dots+\frac{x_{n}-x_{n-1}}{x_{n}-x_{i}}$
$=\bar{h}\left(x_{i};x_{1},\dots,x_{i}\right)+\bar{h}\left(x_{i};x_{i},\dots,x_{n}\right)\ ,$
[…]
A.5 Depth Function of Points Outside the Range of the Dataset
For data points outside the range of the dataset, we have the following theorem.
Theorem A.3.
For any given dataset $D$ with sample size $n>2$ and $x\notin D$, we have
$\bar{h}(x)=\left\{\begin{aligned} \bar{h}(x_{1})\ ,&\quad\text{if }x<x_{1}\ ,\\ \bar{h}(x_{n})\ ,&\quad\text{if }x\geq x_{n}\ ,\\ \ell_{i}(x)\ ,&\quad\text{if }x_{i}\leq x<x_{i+1}\ ,\end{aligned}\right.$
where
$\ell_{i}(x)=\bar{h}(x_{i})+\frac{x-x_{i}}{x_{i+1}-x_{i}}\left(\bar{h}(x_{i+1})-\bar{h}(x_{i})\right)$
is the linear interpolation of $\bar{h}(x_{i})$ and $\bar{h}(x_{i+1})$.
Proof   For $x<x_{1}$ or $x>x_{n}$, evidently, isolating $x$ is equivalent to isolating $x_{1}$ or $x_{n}$, respectively. Thus, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{1};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right],\textrm{for }\forall x<x_{1}$
and
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x_{n};x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right],\textrm{for }\forall x>x_{n}\ .$
We now consider the case where $x\in\left[x_{1},x_{n}\right]$. By Lemma 3.4, it suffices to show that for $x\in\left[x_{i},x_{i+1}\right]$, the depth function $h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)$ is a linear function. For any given $\mathbf{\Theta}=\mathbf{\Theta}_{0}$, we have
$h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}_{0}\right)$
is piecewise-constant. Note that every split that does not isolation $x$ may not change the value of $h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}_{0}\right)$. Therefore, we will focus on the split that isolates $x$. Denote by $\Omega$ be the event that $x$ is eventually isolated by a split in $(x_{i},x)$. Then, we have
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=$ $\ \Pr[\mathbf{\Theta}\in\Omega]~{}\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\mid\Theta\in\Omega\right]$
$+\Pr[\mathbf{\Theta}\notin\Omega]~{}\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\mid\Theta\notin\Omega\right]\ .$
When the split isolating $x$ is in interval $(x_{i},x)$, point $x$ is treated the same as $x_{i+1}$ and thus has the same depth as $x_{i+1}$. Similarly, when the split isolating $x$ is in interval $(x,x_{i+1})$, point $x$ is treated the same as $x_{i}$ and thus has the same depth as $x_{i}$. Therefore, we have the depth of $x$ is a constant given either $\mathbf{\Theta}\in\Omega$ or $\mathbf{\Theta}\notin\Omega$, which implies that
$\mathbb{E}_{\mathbf{\Theta}}\left[h\left(x;x_{1},\dots,x_{n},\mathbf{\Theta}\right)\right]=\Pr[\mathbf{\Theta}\in\Omega]c_{1}+\Pr[\mathbf{\Theta}\notin\Omega]c_{2}\ ,$
where $c_{1}$ and $c_{2}$ are constants. Then it suffices to analyze the probability of event $\Omega$. By the growth mechanism of isolation forest, the split that isolates $x$ is uniformly sampled from the interval $(x_{i},x_{i+1})$, and thus we have
$\Pr[\mathbf{\Theta}\in\Omega]=\frac{x-x_{i}}{x_{i+1}-x_{i}}\quad\textrm{and}\quad\Pr[\mathbf{\Theta}\notin\Omega]=\frac{x_{i+1}-x}{x_{i+1}-x_{i}}\ .$
As the dataset is fixed, $\Pr[\mathbf{\Theta}\in\Omega]$ and $\Pr[\mathbf{\Theta}\notin\Omega]$ are both linear in $x$, which completes the proof.
[…]
Appendix B Further Elaboration of Assumption 4.2
Recall that we assumed $\kappa>\Omega(\sqrt{n+3})$ in Assumption 4.2. Here, we will show that $\kappa>\Omega(\sqrt{n})$ is commonly satisfied in practice. Note that any distribution can be decomposed into a mixture of uniform distributions, and we will focus on the uniform distribution, for which we have the following proposition.
Proposition B.1.
Let $X_{1},\dots,X_{n},n>3$ be i.i.d. random variables from $\mathcal{U}[0,1]$. Then, with probability $1-O(1/n^{1/2})$, we have the following holds
$\kappa\geq\frac{1}{2}\sqrt{n}\ .$
Proof   Let $L=\min_{i}~{}\lvert X_{i+1}-X_{i}\rvert$. We first show that
$\mathbb{E}[L]=1/(n^{2}-1)\ .$
By the symmetry of order statistics, we have
$\Pr[L>t]=n!\Pr[L>t,X_{1}<\dots<X_{n}]\ .$
Observing that $\Pr[L>t,X_{1}<\dots<X_{n}]$ equals the volume of the set
$S=\{(x_{1},\dots,x_{n})\in[0,1]^{n}\mid x_{i}+t\leq x_{i+1},i=1,\dots,n-1\}\ .$
By applying the transformation
$(y_{1},\dots,y_{n})=(x_{1},x_{2}-t,x_{3}-2t,\dots,x_{n}-(n-1)t)\ ,$
which is a volume-preserving transformation, we have a new set
$S^{\prime}=\{(x_{1},\dots,x_{n})\in[0,1-(n-1)t]^{n}\mid x_{i}+t\leq x_{i+1},i=1,\dots,n-1\}\ .$
Again by the symmetry of order statistics, we have
$\Pr[L>l,X_{1}<\dots<X_{n}]$ $=n!~{}\textrm{Vol}(S^{\prime})$
$=n!~{}\frac{1}{n!}(1-(n-1)t)^{n}=(1-(n-1)t)^{n}\ .$
Therefore, we have
[…]
By Markov’s inequality, we have
$\Pr\left[L>\frac{1}{n^{2}-1}+\frac{1}{n^{3/2}}\right]\leq\frac{n^{3/2}}{n^{2}-1}=O(1/n^{1/2})\ .$
Observing that
$\Pr[\max_{i}X_{i}-\min_{i}X_{i}\leq 1/2]\leq\frac{c}{2^{n}}\ ,$
for some $c>0$, implying that
$\Pr\left[U\leq\frac{1}{2n}\right]\leq\frac{c}{2^{n}}\ ,$
Therefore, we have
$\Pr\left[\frac{U}{L}\leq\frac{\frac{1}{2n}}{\frac{1}{(n^{2}-1)}+\frac{1}{n^{3/2}}}\right]\leq O(n^{-\frac{1}{2}}+2^{-n})=O(n^{-\frac{1}{2}})\ ,$
[…]
Appendix C Proofs of Theorems 4.3-4.5
We will prove Theorems about marginal single anomaly here, including Theorem 4.3-4.5.
C.1 Proof of Theorem 4.3
To avoid ambiguity, we denote by $U_{\textrm{ms}}=U$ and $L_{\textrm{ms}}=L$ following the definitions in Definition 4.1. Suppose that $x_{2}-x_{1}>U_{\textrm{ms}}\cdot\kappa$. We will prove by showing that for all $j>1$, the following holds
$\bar{h}(x_{1};x_{1:n})<\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})\leq\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})\leq\bar{h}(x_{j};x_{1:n})\ ,$
where we take $\sup$ and $\inf$ instead of $\max$ and $\min$ because the maximal and the minimal may not exist when
$x_{2}-x_{1}>U_{\textrm{ms}}\cdot\kappa\ .$
Note that
$\bar{h}(x_{1};x_{1:n})<\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})\quad$
and $\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})\leq\bar{h}(x_{j};x_{1:n})$
are trivial. Then it suffices to show that
$\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})\leq\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})\ .$
For convenience, we define
$s_{i}\triangleq x_{i+1}-x_{i},i\leq n-1\ .$
Apparently, assignment to $(x_{1},\dots,x_{n})$ is equivalent to assigning $(s_{1},\dots,s_{n-1})$. From the condition $U_{\textrm{ms}}=\max_{i\geq 2}~{}\lvert x_{i+1}-x_{i}\rvert$ and $x_{2}-x_{1}>U_{\textrm{ms}}\cdot\kappa$, we have
$s_{1}>U_{\textrm{ms}}\cdot\kappa\geq U_{\textrm{ms}}\geq s_{j},~{}\forall j>1\ .$
We assert that $\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})$ is achieved at $j=n$, i.e.,
$\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})=\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ .$
Otherwise, there exists a $j_{0}$ satisfying $1<j_{0}<n$ such that
$\inf_{x_{1:n}}\bar{h}(x_{j_{0}};x_{1:n})=\inf_{j>1,x_{1:n}}\bar{h}(x_{j};x_{1:n})<\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ .$
By Lemma 3.4, we have
$\bar{h}(x_{j_{0}};x_{1:n})=$ $~{}\frac{s_{1}}{s_{1}+\dots+s_{j_{0}-1}}+\frac{s_{2}}{s_{2}+\dots+s_{j_{0}-1}}+\dots+\frac{s_{j_{0}-1}}{s_{j_{0}-1}}$
$~{}+\frac{s_{j_{0}}}{s_{j_{0}}}+\frac{s_{j_{0}+1}}{s_{j_{0}}+s_{j_{0}+1}}+\dots+\frac{s_{n-1}}{s_{j_{0}}+\dots+s_{n-1}}\ .$
Notice that
$\frac{s_{j_{0}}}{s_{j_{0}}}+\frac{s_{j_{0}+1}}{s_{j_{0}}+s_{j_{0}+1}}$ $+\dots+\frac{s_{n-1}}{s_{j_{0}}+\dots+s_{n-1}}$
$>$ $\ \frac{s_{j_{0}}}{s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}+\frac{s_{j_{0}+1}}{s_{j_{0}+1}+s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}$
$+\dots+\frac{s_{n-1}}{s_{n-1}+\dots+s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}$
$=$ $\ \sum_{i=j_{0}}^{n-1}{\frac{s_{i}}{s_{i}+\dots+s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}}\ .$
Therefore, we have
$\bar{h}(x_{j_{0}};x_{1:n})>\sum_{i=1}^{j_{0}-1}{\frac{s_{i}}{s_{i}+\dots+s_{j_{0}-1}}}+\sum_{i=j_{0}}^{n-1}{\frac{s_{i}}{s_{i}+\dots+s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}}\ .$
If we re-assign $(s^{\prime}_{1},\dots,s^{\prime}_{n-1})$ with $(s_{n-1},\dots,s_{j_{0}},s_{1},\dots,s_{j_{0}-1})$ and re-assign the corresponding $(x^{\prime}_{1},\dots,x^{\prime}_{n})$ with $s_{i}^{\prime}$. Then, we have
$\bar{h}(x^{\prime}_{n};x^{\prime}_{1:n})=\sum_{i=1}^{j_{0}-1}{\frac{s_{i}}{s_{i}+\dots+s_{j_{0}-1}}}+\sum_{i=j_{0}}^{n-1}{\frac{s_{i}}{s_{i}+\dots+s_{j_{0}}+s_{1}+\dots+s_{j_{0}-1}}}<\bar{h}(x_{j_{0}};x_{1:n})\ ,$
which conflicts with the condition that $\bar{h}(x_{j_{0}};x_{1:n})$ is the minimal. To this end, it suffices to show that
$\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})<\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ .$
Recall that
$\bar{h}(x_{1};x_{1:n})=\frac{s_{1}}{s_{1}}+\frac{s_{2}}{s_{1}+s_{2}}+\dots+\frac{s_{n-1}}{s_{1}+\dots+s_{n-1}}\triangleq\tilde{h}(x_{1};s_{1},\dots,s_{n-1})\ ,$
which is decreasing with respect to $s_{1}$ and increasing with respect to $s_{n-1}$, implying that
$\ \sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})$
$=$ $\ \sup_{s_{2},\dots,s_{n-2}}\tilde{h}(x_{1};U_{\textrm{ms}}\cdot\kappa,s_{2},\dots,s_{n-2},U_{\textrm{ms}})$
$=$ $\ \sup_{s_{2},\dots,s_{n-2}}\frac{U_{\textrm{ms}}\cdot\kappa}{U_{\textrm{ms}}\cdot\kappa}+\frac{s_{2}}{U_{\textrm{ms}}\cdot\kappa+s_{2}}+\dots+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}\ .$
We now consider the following optimization problem
$\max_{s_{2},\dots,s_{n-2}}$ $\quad\frac{U_{\textrm{ms}}\cdot\kappa}{U_{\textrm{ms}}\cdot\kappa}+\frac{s_{2}}{U_{\textrm{ms}}\cdot\kappa+s_{2}}+\dots+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}\ ,$
s.t. $\quad L_{\textrm{ms}}\leq s_{i}\leq U_{\textrm{ms}},\quad\forall i\ ,$
which has a Lagrangian function as follows:
$L(s_{2},\dots,s_{n-2})=$ $\ \frac{U_{\textrm{ms}}\cdot\kappa}{U_{\textrm{ms}}\cdot\kappa}+\frac{s_{2}}{U_{\textrm{ms}}\cdot\kappa+s_{2}}+\dots+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}$
$\ +\mu_{2}\left(L_{\textrm{ms}}-s_{2}\right)+\dots+\mu_{n-2}\left(L_{\textrm{ms}}-s_{n-2}\right)$
$\ +\mu_{2}^{\prime}\left(s_{2}-U_{\textrm{ms}}\right)+\dots+\mu_{n-2}^{\prime}\left(s_{n-2}-U_{\textrm{ms}}\right)\ .$
The KKT conditions are
$\frac{\partial L}{\partial s_{i}}=&\ \frac{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i-1}+s_{i})^{2}}-\sum_{i^{\prime}=i+1}^{n-2}{\frac{s_{i^{\prime}}}{(U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{i^{\prime}})^{2}}}\\ &\ -\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}-\mu_{i}+\mu_{i}^{\prime}\\ =&\ 0\ ,\end{aligned}$
$\left(L_{\textrm{ms}}-s_{i}\right)\leq 0\ ,\quad\left(s_{i}-U_{\textrm{ms}}\right)\leq 0\ ,\quad\mu_{i}\leq 0,\quad\mu_{i}^{\prime}\leq 0$
$\mu_{i}\left(L_{\textrm{ms}}-s_{i}\right)=0\ ,\quad\mu_{i}^{\prime}\left(s_{i}-U_{\textrm{ms}}\right)=0\ .$
We have the KKT point as follows
$(s_{2},\dots,s_{n-2})=(U_{\textrm{ms}},\dots,U_{\textrm{ms}})\ .$
To verify the KKT point, we first observe that
$\sum_{i^{\prime}=i+1}^{n-2}{\frac{1}{(U_{\textrm{ms}}\cdot\kappa+(i^{\prime}-1)U_{\textrm{ms}})^{2}}}$ $+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+s_{2}+\dots+s_{n-2}+U_{\textrm{ms}}}$
$=\sum_{i^{\prime}=i+1}^{n-1}{\frac{1}{(U_{\textrm{ms}}\cdot\kappa+(i^{\prime}-1)U_{\textrm{ms}})^{2}}}$
$<\sum_{i^{\prime}=i+1}^{n-1}{\left(\frac{1}{U_{\textrm{ms}}\cdot\kappa+(i^{\prime}-2)U_{\textrm{ms}}}-\frac{1}{U_{\textrm{ms}}\cdot\kappa+(i^{\prime}-1)U_{\textrm{ms}}}\right)}$
$=\frac{1}{U_{\textrm{ms}}\cdot\kappa+(i-1)U_{\textrm{ms}}}-\frac{1}{U_{\textrm{ms}}\cdot\kappa+(n-2)\cdot U_{\textrm{ms}}}\ ,$
Note that the following holds under Assumption 4.2:
$\frac{U_{\textrm{ms}}\cdot\kappa+(i-2)U_{\textrm{ms}}}{(U_{\textrm{ms}}\cdot\kappa+(i-1)U_{\textrm{ms}})^{2}}-\frac{U_{\textrm{ms}}}{(U_{\textrm{ms}}\cdot\kappa+(i-1)U_{\textrm{ms}})}+\frac{U_{\textrm{ms}}}{(U_{\textrm{ms}}\cdot\kappa+(n-2)U_{\textrm{ms}})}>0\ ,$
implying that
$\mu_{i}=0\quad\textrm{and}\quad\mu_{i}^{\prime}<0\ ,$
which verifies the KKT conditions. Therefore, we have
$\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})=\sup_{s_{2},\dots,s_{n-1}}\tilde{h}(x_{1};s_{1},s_{2},\dots,s_{n-1})=\tilde{h}(x_{1};U_{\textrm{ms}}\cdot\kappa,\underbrace{U_{\textrm{ms}},\dots,U_{\textrm{ms}}}_{n-2})\ .$
Similarly, we have
$\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})=\inf_{s_{1},\dots,s_{n-1}}\tilde{h}(x_{n};s_{1},s_{2},\dots,s_{n-1})=\tilde{h}(x_{n};\underbrace{L_{\textrm{ms}},\dots,L_{\textrm{ms}}}_{n-2},U_{\textrm{ms}})\ .$
By Lemma 3.4, we have
$\sup_{x_{1:n}}\bar{h}(x_{1};x_{1:n})$ $=\tilde{h}(x_{1};U_{\textrm{ms}}\cdot\kappa,\underbrace{U_{\textrm{ms}},\dots,U_{\textrm{ms}}}_{n-2})$
$=\frac{U_{\textrm{ms}}\cdot\kappa}{U_{\textrm{ms}}\cdot\kappa}+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+U_{\textrm{ms}}}+\dots+\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}\cdot\kappa+U_{\textrm{ms}}+\dots+U_{\textrm{ms}}+U_{\textrm{ms}}}$
$\leq\frac{U_{\textrm{ms}}}{U_{\textrm{ms}}}+\frac{L_{\textrm{ms}}}{L_{\textrm{ms}}+U_{\textrm{ms}}}+\dots+\frac{L_{\textrm{ms}}}{L_{\textrm{ms}}+\dots+L_{\textrm{ms}}+U_{\textrm{ms}}}$
$=\tilde{h}(x_{n};\underbrace{L_{\textrm{ms}},\dots,L_{\textrm{ms}}}_{n-2},U_{\textrm{ms}})$
$\leq\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ ,$
[…]
C.2 Proof of Theorem 4.4
We prove by giving the assignment of $(x_{1},\dots,x_{n})$ such that when $U_{\textrm{ms}}\cdot\kappa>x_{2}-x_{1}>U_{\textrm{ms}}$
$\exists j>1,\quad\bar{h}(x_{1};x_{1:n})\leq\bar{h}(x_{j};x_{1:n})\ .$
We assign $x_{i}$ from the corresponding $s_{i}$ as follows
$(s_{1},\dots,s_{n-1})=(U_{\textrm{ms}}+\epsilon,U_{\textrm{ms}}/2,U_{\textrm{ms}},\dots,U_{\textrm{ms}})\ ,$
where $\epsilon>0$ is relative small than $U_{\textrm{ms}}$. By mathematical induction on $n$, we can show that
$\tilde{h}(x_{1};U_{\textrm{ms}},U_{\textrm{ms}}/2,U_{\textrm{ms}},\dots,U_{\textrm{ms}})$ $=\tilde{h}(x_{1};1,1/2,1,\dots,1)$
$>\tilde{h}(x_{n};1,1/2,1,\dots,1)$
$=\tilde{h}(x_{n};U_{\textrm{ms}},U_{\textrm{ms}}/2,U_{\textrm{ms}},\dots,U_{\textrm{ms}})$
By the continuity of $\tilde{h}$ and $\bar{h}$, there exists an $\epsilon>0$ such that
$\tilde{h}(x_{1};U_{\textrm{ms}}+\epsilon,U_{\textrm{ms}}/2,U_{\textrm{ms}},\dots,U_{\textrm{ms}})>\tilde{h}(x_{n};U_{\textrm{ms}}+\epsilon,U_{\textrm{ms}}/2,U_{\textrm{ms}},\dots,U_{\textrm{ms}})$
[…]
C.3 Proof of Theorem 4.5
Recall that the output of $k$-Nearest Neighbor is
$h_{knn}(x;D)\triangleq\frac{1}{k}\sum_{x^{\prime}\in\mathcal{N}_{k}(x)}\|x-x^{\prime}\|_{1}\ .$
Similar to the proof of Theorem 4.3, we will prove by showing that for all $j>1$, the following holds
$h_{knn}(x_{1};x_{1:n})>\inf_{x_{1:n}}h_{knn}(x_{1};x_{1:n})\geq\sup_{j^{\prime},x_{1:n}}h_{knn}(x_{j^{\prime}};x_{1:n})\geq h_{knn}(x_{j};x_{1:n}),$
of which the direction of inequalities has reversed direction compared to Theorem 4.3. This arises due to the negative correlation between the depth function and the distance to the nearest neighbor. Similarly, we define
$\tilde{h}_{knn}(x_{1};s_{1},\dots,s_{n-1})\triangleq h_{knn}(x_{1};x_{1:n})\ ,$
where $s_{i}\triangleq x_{i+1}-x_{1}$ are the difference between two neighbored points. Then, we have It is not difficult to show that
$\inf_{x_{1:n}}h_{knn}(x_{1};x_{1:n})$ $=\inf_{s_{1:n-1}}\tilde{h}_{knn}(x_{1};s_{1:n-1})$
$=\tilde{h}_{knn}(s_{1}^{\circ},L_{\textrm{ms}},\dots,L_{\textrm{ms}})$
$=\frac{1}{k}\left[s_{1}^{\circ}+s_{1}^{\circ}+L_{\textrm{ms}}+\dots+s_{1}^{\circ}+(k-1)L_{\textrm{ms}}\right]$
$=\frac{1}{k}\left[\frac{k(k+1)s_{1}^{\circ}}{2}+\frac{k(k-1)}{2}L_{\textrm{ms}}\right]$
$=\frac{k+1}{2}s_{1}^{\circ}+\frac{k-1}{2}L_{\textrm{ms}}$
$=\frac{k+1}{2}U_{\textrm{ms}}\ ,$
where $s_{1}^{\circ}=U_{\textrm{ms}}+(k-1)(U_{\textrm{ms}}-L_{\textrm{ms}})/2$. Similarly, we have
$\sup_{x_{1:n}}h_{knn}(x_{n};x_{1:n})$ $=\sup_{s_{1:n-1}}\tilde{h}_{knn}(x_{n};s_{1:n-1})$
$=\tilde{h}_{knn}(x_{n},U_{\textrm{ms}},\dots,U_{\textrm{ms}})$
$=\frac{1}{k}(U_{\textrm{ms}}+2U_{\textrm{ms}}+\dots+kU_{\textrm{ms}})$
$=\frac{k+1}{2}U_{\textrm{ms}}$
$=\inf_{x_{1:n}}h_{knn}(x_{1};x_{1:n})\ .$
[…]
Appendix D Proofs of Theorems 4.6-4.7
We will prove Theorems 4.6-4.7 in this section.
D.1 Proof of Theorem 4.6
Before that, we first define $U_{\textrm{cs}}$ and $L_{\textrm{cs}}$ as follows:
$U_{\textrm{cs}}=\max_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\quad L_{\textrm{cs}}=\min_{j\neq n_{0}/2}\lvert x_{j+1}-x_{j}\rvert\ .$
Similar to the proof of Theorem 4.6, we define
$\tilde{h}(x_{i};s_{1},\dots,s_{n-1})=\bar{h}(x_{i};x_{1:n})\ ,$
where $x_{i}\in D$ is an arbitrary element in $D$. For convenience, we define
$\theta=\min\{x_{m+1}-x_{m},x_{m+2}-x_{m+1}\}\ .$
We will prove the sufficiency and necessity of $\theta=\Theta(\sqrt{n_{0}})$ separately.
Sufficiency: There exists some constant $c_{1}$ such that when $\theta>c_{1}\cdot\sqrt{n_{0}}$, we have
$\forall j,~{}\bar{h}(x_{n_{0}/2+1};x_{1:n})<\min_{j}\bar{h}(x_{j};x_{1:n})\ .$
By the definition of $\tilde{h}$
$\bar{h}(x_{n_{0}/2+1};x_{1:n})$ $=\tilde{h}(x_{n_{0}/2+1};s_{1},\dots,s_{n-1})$
$<2\sup_{s_{1:n_{0}/2}}\tilde{h}(s_{n_{0}/2};s_{1},\dots,s_{n_{0}/2})$
$\leq 2\tilde{h}(s_{n_{0}/2};U_{\textrm{cs}},\dots,U_{\textrm{cs}},\theta)$
$=2\left(\frac{\theta}{\theta}+\frac{U_{\textrm{cs}}}{\theta+U_{\textrm{cs}}}+\frac{U_{\textrm{cs}}}{\theta+2U_{\textrm{cs}}}+\dots+\frac{U_{\textrm{cs}}}{\theta+(n_{0}/2-1)U_{\textrm{cs}}}\right)$
$\leq 2+2\ln\left[1+\frac{n_{0}/2-1}{\theta/U_{\textrm{cs}}}\right]\ .$
Similarly, we have
$\min_{j}\bar{h}(x_{j};x_{1:n})$ $\geq\inf_{j,s_{1:n}}\bar{h}(x_{j};x_{1:n})$
[…]
$\geq\inf_{s_{1:n_{0}/2-1}}\tilde{h}(x_{1};s_{1:n_{0}/2-1})$
$=\frac{U_{\textrm{cs}}}{U_{\textrm{cs}}}+\frac{L_{\textrm{cs}}}{U_{\textrm{cs}}+L_{\textrm{cs}}}+\frac{L_{\textrm{cs}}}{U_{\textrm{cs}}+2L_{\textrm{cs}}}+\dots+\frac{L_{\textrm{cs}}}{U_{\textrm{cs}}+(n_{0}/2-1)L_{\textrm{cs}}}$
$\geq 1+\ln\left[1+\frac{n_{0}/2-1}{U_{\textrm{cs}}/L_{\textrm{cs}}+1}\right]\ .$
Thus, we expect that
$2+2\ln\left[1+\frac{n_{0}/2-1}{\theta/U_{\textrm{cs}}}\right]$ $=\ln\left[e^{2}\left(1+\frac{n_{0}/2-1}{\theta/U_{\textrm{cs}}}\right)^{2}\right]$
$\leq\ln\left[e\left(1+\frac{n_{0}/2-1}{U_{\textrm{cs}}/L_{\textrm{cs}}+1}\right)\right]=1+\ln\left[1+\frac{n_{0}/2-1}{U_{\textrm{cs}}/L_{\textrm{cs}}+1}\right]\ ,$
which holds if $\theta\geq\Omega(\sqrt{n_{0}})$.
Necessity: If $\theta<o(\sqrt{n_{0}})$, there exists some assignment of $x_{1},\dots,x_{n}$ such that
$\exists j,~{}\bar{h}(x_{n_{0}/2+1};x_{1:n})\geq\bar{h}(x_{j};x_{1:n})\ .$
We prove this by giving an example of such an assignment as follows:
$x_{1}=1,\quad x_{2}=2,\quad\dots,\quad x_{n_{0}/2}=n_{0}/2,\quad x_{n_{0}/2+1}=n_{0}/2+\theta,$
$x_{n_{0}/2+2}=n_{0}/2+2\theta,\quad x_{n_{0}/2+3}=n_{0}/2+2\theta+1,\quad\dots,\quad x_{n}=n_{0}/2+2\theta+n_{0}/2\ ,$
or equivalently,
$(s_{1},\dots,s_{n-1})=\left(1,\dots,1,\theta,\theta,1,\dots,1\right)\ .$
Then we have
$\tilde{h}(x_{n_{0}/2+1})$ $=2\left(\frac{\theta}{\theta}+\frac{1}{\theta+1}+\frac{1}{\theta+2}+\dots+\frac{1}{\theta+(n_{0}/2-1)}\right)$
$\geq 2+2\ln\left(1+\frac{n_{0}/2-1}{\theta+1}\right)$
$\geq\ln\left(\Omega(n_{0}^{2}/\theta^{2})\right)$
and
$\tilde{h}(x_{1})=$ $\ 1+\frac{1}{2}+\frac{1}{3}+\dots+\frac{1}{n_{0}/2-1}+\frac{\theta}{n_{0}/2-1+\theta}+$
$\ \frac{\theta}{2\theta+n_{0}/2-1}+\frac{1}{n_{0}/2-1+2\theta+1}+\dots+\frac{1}{n_{0}/2-1+2\theta+n_{0}/2-1}$
$\leq$ $\ 1+\frac{1}{2}+\frac{1}{3}+\dots+\frac{1}{n_{0}/2-1}+2+\ln(2)$
$\leq$ $\ 3+\ln(2)+\ln\left(n_{0}/2-1\right)$
$=$ $\ \ln(O(n_{0}))$
Therefore, if $\theta=o(\sqrt{n_{0}})$, we have
$\tilde{h}(x_{n_{0}/2+1})=\omega(\tilde{h}(x_{1}))\ ,$
implying that when $n_{0}$ is large enough
$\bar{h}(x_{n_{0}/2+1};x_{1:n})>\bar{h}(x_{1};x_{1:n})\ .$
[…]
D.2 Proof of Theorem 4.7
For convenience, we let $k$ be an even number, otherwise, choose $k+1$ instead of $k$ acts similarly. Following the proof of Theorem 4.5, we have
$\min_{s_{1:n-1}}\tilde{h}_{knn}(x_{n_{0}/2+1};s_{1},\dots,s_{n-1})$ $=\tilde{h}_{knn}(x_{n_{0}/2+1};\underbrace{L_{\textrm{cs}},\dots,L_{\textrm{cs}}}_{n_{0}/2-1},\theta,\theta,\underbrace{L_{\textrm{cs}},\dots,L_{\textrm{cs}}}_{n_{0}/2-1})$
$=\frac{2}{k}\left(\theta+\theta+L_{\textrm{cs}}+\dots\theta+(k/2-1)L_{\textrm{cs}}\right)$
$=\frac{2}{k}\left(k\theta/2+\frac{k/2(k/2-1)}{2}L_{\textrm{cs}}\right)$
$=\theta+\frac{k/2-1}{2}L_{\textrm{cs}}\ .$
Similarly, we have
$\max_{s_{1:n-1}}\tilde{h}_{knn}(x_{1};s_{1},\dots,s_{n-1})$ $=\tilde{h}_{knn}(x_{1};\underbrace{U_{\textrm{cs}},\dots,U_{\textrm{cs}},}_{n_{0}/2-1}\theta,\theta,\underbrace{U_{\textrm{cs}},\dots,U_{\textrm{cs}},}_{n_{0}/2-1})$
$=\frac{1}{k}\left(U_{\textrm{cs}}+2U_{\textrm{cs}}\dots+kU_{\textrm{cs}}\right)$
$=\frac{k+1}{2}U_{\textrm{cs}}\ .$
Then, when $\theta>\frac{k+1}{2}U_{\textrm{cs}}-\frac{k/2-1}{2}L_{\textrm{cs}}=\Omega(k)$, we have $\tilde{h}_{knn}(x_{n_{0}/2+1})>\max_{j}\tilde{h}_{knn}(x_{j})$. The necessity can be proven similarly as in Appendix D.1, by giving an example of such an assignment as follows:
$(s_{1},\dots,s_{n-1})=(\underbrace{1,\dots,1}_{n_{0}/2-1},\theta,\theta,\underbrace{1,\dots,1}_{n_{0}/2-1})\ .$
[…]
Appendix E Proofs of Theorems 4.8-4.9
Here, We will prove Theorems 4.8-4.9. We begin with the following definitions.
$U_{\textrm{mg}}=\max_{1\leq j\leq n_{1}-1}\lvert x_{j+1}-x_{j}\rvert,\quad L_{\textrm{mg}}=\min_{n_{1}+1\leq j\leq n_{1}+n_{0}-1}\lvert x_{j+1}-x_{j}\rvert,\quad\textrm{and }\theta=x_{n_{1}+1}-x_{n_{1}}\ .$
As we assume that $n_{1}$ is odd, we have that $n_{1}=2q+1$ for some $q\in N^{+}$.
E.1 Proof of Theorem 4.8
Sufficiency: There exists some constant $c_{1}$ such that when $\theta\geq c_{1}\cdot n_{1}^{2}$, we have
$\max_{1\leq j\leq n_{1}}\bar{h}(x_{j};x_{1:n})\leq\min_{n_{1}+1\leq j\leq n_{1}+n_{0}}\bar{h}(x_{j};x_{1:n})\ .$
Following the analysis in Appendix C.1, we have
$\max_{1\leq j\leq n_{1}}\bar{h}(x_{j};x_{1:n})=\bar{h}(x_{q+1};x_{1:n})\quad\textrm{and}\quad\min_{2q+2\leq j\leq 2q+1+n_{0}}\bar{h}(x_{j};x_{1:n})=\bar{h}(x_{n};x_{1:n})\ .$
Then, it suffices to show that
$\sup_{x_{1:n}}\bar{h}(x_{q+1};x_{1:n})\leq\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ .$
Similar to the analysis in Appendix C.1, we have
$\sup_{x_{1:n}}\bar{h}(x_{2q+1};x_{1:n})$ $=\sup_{s_{1:q}}\tilde{h}(x_{2q+1};s_{1:q})+\sup_{s_{q+1:2q+n_{0}}}\tilde{h}(x_{2q+1};s_{q+1:2q+n_{0}})$
$=\tilde{h}(x_{q+1};\underbrace{U_{\textrm{mg}},\dots,U_{\textrm{mg}}}_{q-1},L_{\textrm{mg}},L_{\textrm{mg}},\underbrace{U_{\textrm{mg}},\dots,U_{\textrm{mg}}}_{q-1},\theta,\underbrace{U_{\textrm{mg}},\dots,U_{\textrm{mg}}}_{n_{0}-2})$
$=2\left(\frac{L_{\textrm{mg}}}{L_{\textrm{mg}}}+\frac{U_{\textrm{mg}}}{L_{\textrm{mg}}+U_{\textrm{mg}}}+\dots+\frac{U_{\textrm{mg}}}{L_{\textrm{mg}}+(q-1)U_{\textrm{mg}}}\right)+$
$\quad\ \frac{\theta}{2L_{\textrm{mg}}+2(q-1)U_{\textrm{mg}}+\theta}+$
$\quad\ \frac{U_{\textrm{mg}}}{2L_{\textrm{mg}}+2(q-1)U_{\textrm{mg}}+\theta+U_{\textrm{mg}}}+$
[…]
$\quad\ \frac{U_{\textrm{mg}}}{2L_{\textrm{mg}}+2(q-1)U_{\textrm{mg}}+\theta+(n_{0}-2)U_{\textrm{mg}}}$
$\leq\ln O\left(q^{2}n_{0}/(q+\theta)\right)\ .$
and
$\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})$ $=\inf_{s_{1:n-1}}\tilde{h}(x_{n};s_{1:n-1})$
$=\tilde{h}(x_{n};\underbrace{L_{\textrm{mg}},\dots,L_{\textrm{mg}}}_{2q},\theta,\underbrace{L_{\textrm{mg}},\dots,L_{\textrm{mg}}}_{n_{0}-2},U_{\textrm{mg}})$
$=\frac{U_{\textrm{mg}}}{U_{\textrm{mg}}}+\frac{L_{\textrm{mg}}}{U_{\textrm{mg}}+L_{\textrm{mg}}}+\dots+\frac{L_{\textrm{mg}}}{U_{\textrm{mg}}+(n_{0}-2)U_{\textrm{mg}}}+*\quad$
$\geq\ln\left(\Omega\left(n_{0}\right)\right)\ ,$
where * is some positive quantity. Therefore, $\theta=\Omega(q^{2})=\Omega(n_{1}^{2})$ suffices to ensure that
$\sup_{x_{1:n}}\bar{h}(x_{q+1};x_{1:n})\leq\inf_{x_{1:n}}\bar{h}(x_{n};x_{1:n})\ ,$
implying that
$\max_{1\leq j\leq n_{1}}\bar{h}(x_{j};x_{1:n})\leq\min_{n_{1}+1\leq j\leq n_{1}+n_{0}}\bar{h}(x_{j};x_{1:n})\ .$
Neccessity: We will show that if $\theta=o(n_{1}^{2})$, then there exists some assignment of $x_{1:n}$ such that
$\max_{1\leq j\leq n_{1}}\bar{h}(x_{j};x_{1:n})>\min_{n_{1}+1\leq j\leq n_{1}+n_{0}}\bar{h}(x_{j};x_{1:n})\ .$
We similarly give the assignment of $x_{1:n}$ as follows:
$(s_{1},\dots,s_{n_{1}-1})=(\underbrace{1,\dots,1}_{2q},\theta,\underbrace{1,\dots,1}_{n_{0}-1})\ .$
Then, we have
$\tilde{h}(x_{q+1};s_{1:n-1})$ $=2\left(1+\frac{1}{2}+\dots\frac{1}{q}\right)+\frac{\theta}{2q+\theta}+\frac{1}{2q+\theta+1}+\dots+\frac{1}{2q+\theta+n_{0}-2}$
$\geq\ln\left(\Omega\left(q^{2}n_{0}/(q+\theta)\right)\right)$
and
$\tilde{h}(x_{n};s_{1:n-1})=$ $\ 1+\frac{1}{2}+\frac{1}{3}+\dots+\frac{1}{n_{0}-2}+\frac{\theta}{n_{0}-2+\theta+1}$
$+\frac{1}{n_{0}-2+\theta+2}+\dots+\frac{1}{n_{0}-2+\theta+2q}\leq\ln O\left(n_{0}\right)\ .$
Thus, if $\theta=o(n_{1}^{2})$, then
$\max_{1\leq j\leq n_{1}}\bar{h}(x_{j};x_{1:n})>\min_{n_{1}+1\leq j\leq n_{1}+n_{0}}\bar{h}(x_{j};x_{1:n})\ .$
[…]
E.2 Proof of Theorem 4.9
Sufficiency: There exists some constant $c_{1}$ such that when $\theta\geq c_{1}k$, we have
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})\geq\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})\ .$
Observing that
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})$ $=\inf_{1\leq j\leq n_{1},s_{1:n-1}}\tilde{h}(x_{j};s_{1:n-1})$
$=\inf_{1\leq j\leq n_{1},s_{1:n_{1}-1}}\tilde{h}(x_{j};s_{1:n_{1}-1},\theta,\underbrace{L_{\textrm{mg}},\dots,L_{\textrm{mg}}}_{n_{0}-1})$
$\geq\frac{1}{k}\left[(k-n_{1}-2)\theta+(k-n_{1}-2)(k-n_{1}-3)L_{\textrm{mg}}/2\right]$
$=\Omega(\theta+k)$
and
$\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})$ $=\sup_{n_{1}+1\leq j\leq n_{1}+n_{0},s_{1:n-1}}\tilde{h}(x_{j};s_{1:n-1})$
$=\sup_{n_{1}+1\leq j\leq n_{1}+n_{0},s_{1:n_{1}-1}}\tilde{h}(x_{n};s_{1:n_{1}-1},\theta,\underbrace{U_{\textrm{mg}},\dots,U_{\textrm{mg}}}_{n_{0}-1})$
$\leq\frac{1}{k}\left(U_{\textrm{mg}}+2U_{\textrm{mg}}+\dots+kU_{\textrm{mg}}\right)$
$=O(k)\ .$
Therefore, $\theta=\Omega(k)$ suffices to ensure that
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})\geq\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})\ .$
Neccessity: We will show that if $\theta=o(k)$, then there exists some assignment of $x_{1:n}$ such that
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})<\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})\ .$
We similarly give the assignment of $x_{1:n}$ corresponding to $s_{1:n-1}$ as follows:
$(s_{1},\dots,s_{n-1})=(\underbrace{L_{\textrm{mg}},\dots,L_{\textrm{mg}}}_{n_{1}-1},\theta,\underbrace{L_{\textrm{mg}},\dots,L_{\textrm{mg}}}_{k-n_{1}-2},\underbrace{U_{\textrm{mg}},\dots,U_{\textrm{mg}}}_{n-k+1})\ .$
Then, we have
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})$ $\leq\frac{1}{k}\Big{(}L_{\textrm{mg}}+2L_{\textrm{mg}}+\dots+(n_{1}-1)L_{\textrm{mg}}+\theta+(\theta+L_{\textrm{mg}})+\dots+(\theta+(k-n_{1})L_{\textrm{mg}})\Big{)}$
$\leq\left[\frac{n_{1}(n_{1}-1)}{2}+\frac{(k-n_{1})(k-n_{1}-1)}{2}\right]\frac{L_{\textrm{mg}}}{k}+\theta$
and
$\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})$ $\geq\frac{1}{k}(1+2+\dots+k)U_{\textrm{mg}}\ .$
Thus, we have
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})-\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})\leq-\Omega(k)+\theta\ .$
Once $\theta=o(k)$, we have
$\min_{1\leq j\leq n_{1}}h_{knn}(x_{j};x_{1:n})-\max_{n_{1}+1\leq j\leq n_{1}+n_{0}}h_{knn}(x_{j};x_{1:n})\leq-\Omega(k)<0\ .$
[…]